OpenStudy (anonymous):

Write an equation to represent the relationship, and then solve the equation. 25 more than 3 times a number is the same as 9 times the number minus two

OpenStudy (accidentalaichan):

25 + 3x = 9x - 2. Solve

OpenStudy (accidentalaichan):

If you need help, just ask

OpenStudy (accidentalaichan):

And remember the instructions: 1. Combine like terms 2. rewrite equation 3. Use inverse operations

OpenStudy (anonymous):

is the answer x=4.5?

OpenStudy (accidentalaichan):

Let's see here 25 + 3x = 9x - 2 25 = 6x - 2 2 + 25 = 27 27 / 6 = 4.5 4.5 = x So yes, you are correct

OpenStudy (anonymous):

Yay, thanks for the help.

OpenStudy (accidentalaichan):

Np